ABC Company needs one additional worker for two-year project. It could hire an employee for a $65,000 annual salary. Assume that ABC does not provide health insurance to employees. Instead of hiring a new employee, ABC could engage an independent contractor for $78,000. If ABC's marginal tax rate is 35%, which option minimizes the after-tax cost of obtaining the worker?
Answer #1

Annual salary if the employee is hired= 65000

If an independent contractor is engaged = 78000

tax implication on contractor =78000*35%=27300

net cost on contract= 50700 (78000-27300)

therefore it is preferential to have an independent contractor